OpenStudy (anonymous):

The function below models the correlation between the number of hours a plant is kept in sunlight (x) and the height (y), in mm, to which it grows: y = 2 + 4x What does the y-intercept of this function represent? The original height of the plant was four mm. The original height of the plant was two mm. The height of the plant increases by two mm for every hour of sunlight it receives. The height of the plant increases by four mm for every hour of sunlight it receives.

OpenStudy (ybarrap):

Which term is the y-intercept?

OpenStudy (ybarrap):

When is y equal to it? This tells you the meaning of the y-intercept.

OpenStudy (ybarrap):

y=mx+b is the general form of a linear equation m is the slope b is the y-intercept when x = 0 what is y?

OpenStudy (ybarrap):

What does x=0 physically mean?

OpenStudy (anonymous):

Not really sure :/ what would it be?

OpenStudy (anonymous):

@ybarrap

OpenStudy (ybarrap):

x represents the amount of sunlight. So if x = 0 that means, what?

OpenStudy (ybarrap):

if you have 0 sunlight, what does that mean?

OpenStudy (anonymous):

it means you have no sunlight

OpenStudy (ybarrap):

It means it's dark :)

OpenStudy (ybarrap):

So, no light, means that plant can't grow. y = 2 + 4x When no sunlight is present y=2 No growth Plant stays the same height.

OpenStudy (anonymous):

So which option would you say is correct?

OpenStudy (anonymous):

@ybarrap

OpenStudy (ybarrap):

Plant stays the same height. If x=0 the plant doesn't grow, it's original length was y=2 (mm) so it stays at that height.

OpenStudy (anonymous):

Alright thank you!